# ReservoirTTA: Prolonged Test-time Adaptation for Evolving and Recurring Domains

### v1

@@ Abstract
[…]
@@ 1 Introduction
Deep networks have achieved state-of-the-art performance across many tasks, but their reliability degrades when test-time data deviates from the training distribution. Real-world deployment scenarios such as autonomous driving or surveillance often involve dynamic shifts caused by changing weather, sensor degradation, or environmental variation. These settings call for robust test-time adaptation (TTA) methods [36, 23, 43, 19, 32, 33] that enable pre-trained models to adapt on-the-fly, ideally over prolonged periods, without catastrophic forgetting or model collapse. Most existing TTA methods, e.g., efficient TTA (ETA) [24], assume each domain appears only once in the test stream. In real-world long-term deployments, however, domain conditions often recur. As Figure 1 (left) shows, visual distributions may shift and later reappear. Empirically, this recurring behavior destabilizes ETA [24], which lacks explicit regularization, whereas anti-forgetting TTA (EATA) [24] maintains greater long-term stability by constraining parameter drift. Even so, when domains re-emerge, these regularized methods remain vulnerable to catastrophic forgetting, as illustrated in Figure 1 (right).
To address this challenge, we clarify two related yet distinct axes of variation—style and domain—and revisit how their boundaries are detected. Prior work typically treats domains as discrete, source-annotated groups (e.g., different sensors or collection conditions). Styles and domains may not map one-to-one: a single domain can exhibit multiple styles due to intra-domain variability, and distinct domains can share similar style signatures (for example, zoom blur and defocus blur yield nearly identical style statistics). Nevertheless, in our framework, a domain is represented by a group of similar styles, where style is computed as the set of low-level appearance attributes captured via channel-wise feature statistics from early layers of a pre-trained VGG network [30]. This allows us to leverage the style embeddings to detect new domains or recognize the recurring ones via online clustering without over-fragmentation—an issue observed in the recent methods lacking robust style representations (see  Figure 16)  [43, 2]. Based on the style of the current test samples, we route them to the corresponding domain model in our reservoir, where specialized adaptation is performed.
In summary, we introduce ReservoirTTA, which maintains a pool of domain-specialized models, adapts each one independently with its corresponding TTA updates, and combines their parameters in a weighted ensemble for final predictions. Our contributions are as follows:
@@
  • Theoretical insights into long-term stability. We provide theoretical bounds showing how parameter regularization curbs collapse in single-domain TTA—clarifying the stability of methods such as EATA [24]—and motivate the modular design of ReservoirTTA for maintaining stability under recurring domain shifts.
@@ 2 Related Work
Prior works on prolonged, recurring test-time adaptation fall into two main strands. Continual TTA methods [38, 41, 31, 31] continually update a single model to track evolving domains but suffer from drift and forgetting when domains recur, while robust/persistent TTA methods [27, 11, 24, 21] employ techniques such as variance constraints or periodic resets to preserve stability yet lack efficient re-adaptation to previously seen shifts. For a concise survey of representative algorithms and their limitations in recurring and evolving settings, see Appendix A.
@@ 3.2 Test-Time Adaptation Trajectory
The above analysis reveals why conventional TTA approaches often suffer from model collapse during prolonged adaptation: as the number of update steps increases, so does the variance of model parameters, eventually causing them to drift beyond the stability radius. The proof of Theorem 1 is provided in Remark 1. Next, we examine strategies to mitigate this variance growth.
@@
~$\nabla\tilde{\mathcal{L}}_{\text{TTA}}(\theta,\mathbf{x})=S(\mathbf{x})\,\nabla\mathcal{L}_{\text{TTA}}(\theta,\mathbf{x}),$ (3)
thereby reducing the effective gradient variance to $V_{\text{eff}}$ with $V_{\text{eff}}<\bar{V}$, but maintains linear dependence on $t$.
@@
Parameter Drift in Recurring Continual TTA. Even with these variance control strategies (e.g., sample filtering and weight ensembling), a single adapting model remains vulnerable to parameter drift when the shift between domain-optimal parameters exceeds the stability radius, i.e., $\|\theta^{*}_{\text{Task}_{i}}-\theta^{*}_{\text{Task}_{i+1}}\|>\beta_{i+1},\ \text{for some }i.$ As illustrated in Figure 4 of Appendix B.2, such shifts cause the model’s parameter trajectory $\{\theta_{t}\}$ to deviate from the optimal region for a given domain, leading to catastrophic forgetting and negative transfer.
@@ 4 Methodology
In this section, we introduce our framework, ReservoirTTA, for prolonged TTA in environments with recurring and evolving domains. As illustrated in Figure 2, our framework comprises four stages: (1) Style Characterization and Domain Identification—leveraging style-based online clustering to determine the domain characteristics of the current test batch; (2) Model Reservoir Initialization—allocating a new domain-specific model when a novel domain is detected in the domain identification step; (3) Model Reservoir Adaptation—selectively updating only the model associated with the current domain using state-of-the-art TTA techniques; and (4) Model Prediction—Predictions are then obtained via the ensemble’s parameters from all domain-specific models (see pseudocode in Appendix C).
@@ 4.1 Style Characterization and Domain Identification
Style Characterization. To discriminate between domains at test time, we capture the style characteristics of the current test images using feature statistics derived from a pre-trained deep network. Inspired by prior work on style characterization [8, 14, 37] using VGG-19 [30], we compute the batch-wise log-variance of the intermediate features from $L$ early convolutional layers of the VGG-19 $g_{\text{style}}$. Formally, given a batch $\mathbf{B}_{t}\sim\mathcal{D}_{t}$ and feature maps $\{\mathbf{z}_{1},\dots,\mathbf{z}_{L}\}$ with $\mathbf{z}_{l}\in\mathbb{R}^{b\times h_{l}\times w_{l}\times c_{l}}$, the style vector for each layer is computed as $\mathbf{s}_{l}(\mathbf{B}_{t})=\texttt{logvar}(\mathbf{z}_{l})$ where logvar computes the natural logarithm of the variance over the batch, height, and width dimensions. The overall style representation is formed by concatenating the vectors: $\mathbf{s}_{t}=\bigl{[}\mathbf{s}_{1}(\mathbf{B}_{t}),\dots,\mathbf{s}_{L}(\mathbf{B}_{t})\bigr{]}\in\mathbb{R}^{d}.$ We study different configurations of VGG-19 layers and style representations in Appendix D.
@@
To adapt the centroids over time, we optimize a mutual information loss $\mathcal{L}_{\text{MI}}$ between past style vectors and the current centroids. Storing all past styles is infeasible, so we maintain a fixed-size Style Reservoir $\mathbf{R}_{t}:=[\mathbf{s}_{t_{1}},\dots,\mathbf{s}_{t_{M}}]$, updated via Reservoir Sampling [35, 3] to approximate uniform sampling over all previously seen $\mathbf{s}_{t}$. At each step, $\mathbf{s}_{t}$ is added to $\mathbf{R}_{t-1}$ if $t\leq M$; otherwise, it replaces a randomly selected vector with probability $\frac{M}{t}$, ensuring unbiased coverage. We define the soft-assignment for each style vector in the style reservoir to the centroids as follows:
@@
Here, $\mathcal{L}_{\text{ent}}(\mathbf{Q})=-\frac{1}{M}\sum_{i=1}^{M}\sum_{j=1}^{K}q_{ij}\log q_{ij}$ encourages confident assignments of style vectors in $\mathbf{R}_{t}$ to the nearest style centroids, while $\mathcal{L}_{\text{cm}}(\mathbf{Q})=\sum_{j=1}^{K}\bar{q}_{j}\log\bar{q}_{j}$ (with $\bar{q}_{j}=\frac{1}{M}\sum_{i=1}^{M}q_{ij}$) acts to prevent centroid collapse by promoting assignment diversity. Appendix E presents the sensitivity of all introduced hyperparameters, alternative online clustering strategies, and the style reservoir update.
@@ 4.2 Model Reservoir: Initialization, Adaptation, and Prediction
Building on the theoretical analysis of parameter drift (Section 3.2), we maintain a Model Reservoir comprising $K_{t}$ domain-specialized models $\{\theta^{1}_{t},\dots,\theta^{K_{t}}_{t}\}$, with one model per discovered domain up to time $t$. To ensure computational efficiency, we store only the trainable parameters of each model, rather than full model instances. The total number of domains is bounded by a fixed constant $K^{\text{max}}$ to prevent memory exhaustion. At initialization, the reservoir contains a single model $\{\theta^{1}_{0}\}$ corresponding to the source domain. When a new domain is detected, we instantiate a new model by cloning the parameters of an existing reservoir model that yields the most confident and diverse predictions on the current test batch $\mathbf{B}_{t}$. This is formalized using mutual information as follows:
@@
where $f_{\theta}(\mathbf{B}_{t})\in\mathbb{R}^{b\times|Y|}$ is the softmax output of the model. This criterion favors models whose predictions are simultaneously confident and diverse, reducing the risk of collapse when adapting to a novel domain (see alternative initialization in Appendix F).
@@ 5 Experiments
Baselines. We evaluate our method against several state-of-the-art TTA baselines (for more details, see the Appendix G). For single-target TTA, we compare with TENT [36]. For continual TTA, we consider CoTTA [38] (using the affine-parameter variant, CoTTA*), RoTTA [41], ETA [24], and SAR [25]. To assess long-term stability, we include persistent TTA methods such as RDumb [27], PeTTA [11], EATA [24] (with Fisher-based regularization), and ROID [21] (reported as ROID*—a version omitting the augmentation consistency loss). For ViT-based models, we also compare with domain-disentanglement approaches like CoLA [2] and DPCore [43]. In segmentation experiments, we evaluate segmentation variants of TENT, CoTTA*, and BECoTTA [18]. For fair comparison, all methods update only the backbone’s affine parameters—except when we compare ReservoirTTA and BECoTTA (using LoRA) and DPcore (using visual prompts).
@@
Implementation Details. All methods are re-implemented in PyTorch [26] within a unified TTA repository [21] for fair comparison, using pre-trained source models from RobustBench [10]. See Appendix H for further implementation details.
@@ 5.1 Main Results
Classification. Table 1 highlights the limitations of existing TTA methods when repeatedly adapted under recurring domain shifts and prolonged testing with three CNN-based backbones. (1) Single-target TTA: Methods like TENT initially adapt well but suffer from severe error accumulation over multiple cycles—for example, on CIFAR-10-C, error rises from 19.3% at visit 1 to 87.8% at visit 20, with similar trends on CIFAR-100-C and ImageNet-C. Plugging in ReservoirTTA significantly improves performance, though error accumulation still occurs, highlighting the need for both variance control modules from our theoretical analysis. (2) Continual TTA: Approaches like CoTTA and RoTTA mitigate short-term instability yet still accumulate errors with repeated adaptation. ReservoirTTA decouples domain-specific adaptation to prevent parameter drift—reducing, for instance, ETA’s CIFAR-10-C error from 30.9% to 16.4% at the 20th visit (Recurring CSC) and similarly improving ImageNet-C performance. (3) Persistent TTA: Although methods such as EATA and ROID are designed to prevent catastrophic forgetting, they exhibit limited re-adaptation (Figure 1), with minimal improvements over repeated visits. ReservoirTTA overcomes this by enabling controlled adaptation without excessive re-learning—reducing EATA’s ImageNet-C error from 55.9% to 51.0% in Recurring CSC and ROID’s from 55.5% to 52.1%.
As shown in Table 2, using a ViT-B-16 backbone on ImageNet-C further underscores ReservoirTTA’s advantages in recurrent continual TTA. Domain-disentangled methods like DPCore improve early adaptation but lack long-term stability. For instance, in Recurring CSC, CoLA reduces ETA’s 20th-visit error to 33.8%, whereas ReservoirTTA lowers it further to 31.9%. In CCC, CoLA stabilizes ETA at 40.2%, whereas ReservoirTTA achieves 38.5%. DPCore also struggles: in CCC, its error rises from 42.2% to 43.1%, whereas combining ReservoirTTA with the same prompt-tuning approach (VPT) [13]—used by DPCore for fair comparison—keeps the error at 39.2%. This gap shows that the domain identification mechanism in ReservoirTTA outperforms those in CoLA and DPCore by avoiding gradual degradation and accelerating performance improvement. As shown in Appendix D, our style features form higher-quality clusters than the ViT-based features used in CoLA and DPCore. Moreover, our model discovery mechanism is less sensitive to test batch order and estimates the number of domains more accurately (see Appendix I), resulting in fewer trainable parameters compared to DPCore and COLA. Full results on the 20 recurrences and CNN backbones (group norm) can be found in Tables 920 in Appendix K.
@@
[…]
@@
[…]
@@
Recurring visit $\xrightarrow{\hskip 51.21504pt}$ Recurring visit $\xrightarrow{\hskip 51.21504pt}$ Recurring visit $\xrightarrow{\hskip 51.21504pt}$ Recurring visit $\xrightarrow{\hskip 51.21504pt}$
@@
Figure 3: Sensitivity to the maximum number of models in the reservoir $K^{\text{max}}$ on CIFAR100-C under recurring CSC and CDC settings.

### v3

@@ Abstract
This paper introduces ReservoirTTA, a novel plug–in framework designed for prolonged test–time adaptation (TTA) in scenarios where the test domain continuously shifts over time, including cases where domains recur or evolve gradually. At its core, ReservoirTTA maintains a reservoir of domain-specialized models—an adaptive test-time model ensemble—that both detects new domains via online clustering over style features of incoming samples and routes each sample to the appropriate specialized model, and thereby enables domain-specific adaptation. This multi-model strategy overcomes key limitations of single model adaptation, such as catastrophic forgetting, inter-domain interference, and error accumulation, ensuring robust and stable performance on sustained non-stationary test distributions. Our theoretical analysis reveals key components that bound parameter variance and prevent model collapse, while our plug–in TTA module mitigates catastrophic forgetting of previously encountered domains. Extensive experiments on scene-level corruption benchmarks (ImageNet-C, CIFAR-10/100-C), object-level style shifts (DomainNet-126, PACS), and semantic segmentation (Cityscapes→ACDC) — covering recurring and continuously evolving domain shifts — show that ReservoirTTA substantially improves adaptation accuracy and maintains stable performance across prolonged, recurring shifts, outperforming state-of-the-art methods. Our code is publicly available at https://github.com/LTS5/ReservoirTTA.
@@ 1 Introduction
Deep networks have achieved state-of-the-art performance across many tasks, but their reliability degrades when test-time data deviates from the training distribution. Real-world deployment scenarios such as autonomous driving or surveillance often involve dynamic shifts caused by changing weather, sensor degradation, or environmental variation. These settings call for robust test-time adaptation (TTA) methods [25, 46, 20, 39, 36] that enable pre-trained models to adapt on-the-fly, ideally over prolonged periods, without catastrophic forgetting or model collapse. Most existing TTA methods, e.g., efficient TTA (ETA) [26], assume each domain appears only once in the test stream. In real‐world long‐term deployments, however, domain conditions often recur. As Figure 1 (left) shows, visual distributions may shift and later reappear. Empirically, this recurring behavior destabilizes ETA [26], which lacks explicit regularization, whereas anti‐forgetting TTA (EATA) [26] maintains greater long‐term stability by constraining parameter drift. Even so, when domains re‐emerge, these regularized methods remain vulnerable to catastrophic forgetting, as illustrated in Figure 1 (right).
To address this challenge, we clarify two related yet distinct axes of variation—style and domain—and revisit how their boundaries are detected. Prior work typically treats domains as discrete, source‐annotated groups (e.g., different sensors or collection conditions). Styles and domains may not map one‐to‐one: a single domain can exhibit multiple styles due to intra‐domain variability, and distinct domains can share similar style signatures (for example, zoom blur and defocus blur yield nearly identical style statistics). Nevertheless, in our framework, a domain is represented by a group of similar styles, where style is computed as the set of low‐level appearance attributes captured via channel‐wise feature statistics from early layers of a pre‐trained VGG network [34]. This allows us to leverage the style embeddings to detect new domains or recognize the recurring ones via online clustering without over-fragmentation—an issue observed in the recent methods lacking robust style representations (see  Figure˜17)  [46, 3]. Based on the style of the current test samples, we route them to the corresponding domain model in our reservoir, where specialized adaptation is performed.
In summary, we introduce ReservoirTTA, which maintains a pool of domain‐specialized models, adapts each one independently with its corresponding TTA updates, and combines their parameters in a weighted ensemble for final predictions. Our contributions are as follows:
@@
  • Theoretical insights into long-term stability. We provide theoretical bounds showing how parameter regularization curbs collapse in single‐domain TTA—clarifying the stability of methods such as EATA [26]—and motivate the modular design of ReservoirTTA for maintaining stability under recurring domain shifts.
@@ 2 Related Work
Prior works on prolonged, recurring test‐time adaptation fall into two main strands. Continual TTA methods [41, 44, 35] continually update a single model to track evolving domains but suffer from drift and forgetting when domains recur, while robust/persistent TTA methods [30, 12, 26, 23] employ techniques such as variance constraints or periodic resets to preserve stability yet lack efficient re‐adaptation to previously seen shifts. For a concise survey of representative algorithms and their limitations in recurring and evolving settings, see Appendix˜A.
@@ 3.2 Test-Time Adaptation Trajectory
The above analysis reveals why conventional TTA approaches often suffer from model collapse during prolonged adaptation: as the number of update steps increases, so does the variance of model parameters, eventually causing them to drift beyond the stability radius. The proof of Theorem˜1 is provided in Remark˜1. Next, we examine strategies to mitigate this variance growth.
@@
$\nabla\tilde{\mathcal{L}}_{\text{TTA}}(\theta,\mathbf{x})=S(\mathbf{x})\,\nabla\mathcal{L}_{\text{TTA}}(\theta,\mathbf{x}),$ (3)
thereby reducing effective gradient variance to $V_{\text{eff}}<\bar{V}$ while preserving linear dependence on $t$.
@@
Parameter Drift in Recurring Continual TTA. Even with these variance control strategies (e.g., sample filtering and weight ensembling), a single adapting model remains vulnerable to parameter drift when the shift between domain-optimal parameters exceeds the stability radius, i.e., $\|\theta^{*}_{\text{Task}_{i}}-\theta^{*}_{\text{Task}_{i+1}}\|>\beta_{i+1},\ \text{for some }i.$ As illustrated in Figure˜5 of Appendix B.2, such shifts cause the model’s parameter trajectory $\{\theta_{t}\}$ to deviate from the optimal region for a given domain, leading to catastrophic forgetting and negative transfer.
@@ 4 Methodology
In this section, we introduce our framework, ReservoirTTA, for prolonged TTA in environments with recurring and evolving domains. As illustrated in Figure˜2, our framework comprises four stages: (1) Style Characterization and Domain Identification—leveraging style-based online clustering to determine the domain characteristics of the current test batch; (2) Model Reservoir Initialization—allocating a new domain-specific model when a novel domain is detected in the domain identification step; (3) Model Reservoir Adaptation—selectively updating only the model associated with the current domain using state-of-the-art TTA techniques; and (4) Model Prediction—Predictions are then obtained via the ensemble’s parameters from all domain-specific models (see pseudocode in Appendix˜C).
@@ 4.1 Style Characterization and Domain Identification
Style Characterization. To distinguish domains at test time, following prior style-characterization work [9, 15, 40, 34], we encode image style as batch-wise log-variances from the first $L$ layers of a frozen, ImageNet-trained VGG-19 $g_{\text{style}}$. Given a batch $\mathbf{B}_{t}\sim\mathcal{D}_{t}$ and feature maps $\{\mathbf{z}_{1},\dots,\mathbf{z}_{L}\}$ with $\mathbf{z}_{l}\in\mathbb{R}^{b\times h_{l}\times w_{l}\times c_{l}}$, the style vector for layer $l$ is computed as $\mathbf{s}_{l}(\mathbf{B}_{t})=\texttt{logvar}(\mathbf{z}_{l})$ where logvar computes the natural logarithm of the variance over the batch, height, and width. The overall style descriptor is formed by concatenating these vectors: $\mathbf{s}_{t}=\bigl[\mathbf{s}_{1}(\mathbf{B}_{t}),\dots,\mathbf{s}_{L}(\mathbf{B}_{t})\bigr]\in\mathbb{R}^{d}.$ This architecture-agnostic statistic captures robust texture cues while remaining independent of the source model; Appendix˜D analyzes VGG configurations and shows that VGG-based style vectors yield more stable, higher-quality style clustering than source-model and ViT-based alternatives.
@@
To adapt the centroids over time, we optimize a mutual information loss $\mathcal{L}_{\text{MI}}$ between past style vectors and the current centroids. Storing all past styles is infeasible, so we maintain a fixed-size Style Reservoir $\mathbf{R}_{t}:=[\mathbf{s}_{t_{1}},\dots,\mathbf{s}_{t_{M}}]$, updated via Reservoir Sampling [38, 4] to approximate uniform sampling over all previously seen $\mathbf{s}_{t}$. At each step, $\mathbf{s}_{t}$ is added to $\mathbf{R}_{t-1}$ if $t\leq M$; otherwise, it replaces a randomly selected vector with probability $\frac{M}{t}$, ensuring unbiased coverage.
We define the soft-assignment for each style vector in the style reservoir to the centroids as follows:
@@
Here, $\mathcal{L}_{\text{ent}}(\mathbf{Q})=-\frac{1}{M}\sum_{i=1}^{M}\sum_{j=1}^{K}q_{ij}\log q_{ij}$ encourages confident assignments of style vectors in $\mathbf{R}_{t}$ to the nearest style centroids, while $\mathcal{L}_{\text{cm}}(\mathbf{Q})=\sum_{j=1}^{K}\bar{q}_{j}\log\bar{q}_{j}$ (with $\bar{q}_{j}=\frac{1}{M}\sum_{i=1}^{M}q_{ij}$) acts to prevent centroid collapse by promoting assignment diversity. Appendix˜E presents the sensitivity of all introduced hyperparameters, alternative online clustering strategies, and alternative distance metrics, including cosine similarity and Kullback–Leibler (KL) divergence, as well as the style reservoir update.
@@ 4.2 Model Reservoir: Initialization, Adaptation, and Prediction
Building on the theoretical analysis of parameter drift (Section˜3.2), we maintain a Model Reservoir comprising $K_{t}$ domain-specialized models $\{\theta^{1}_{t},\dots,\theta^{K_{t}}_{t}\}$, with one model per discovered domain up to time $t$. To ensure computational efficiency, we store only the trainable parameters of each model, rather than full model instances. The total number of domains is bounded by a fixed constant $K^{\text{max}}$ to prevent memory exhaustion. At initialization, the reservoir contains a single model $\{\theta^{1}_{0}\}$ corresponding to the source domain. When a new domain is detected, we instantiate a new model by cloning the parameters of an existing reservoir model that yields the most confident and diverse predictions on the current test batch $\mathbf{B}_{t}$. This is formalized using mutual information as follows:
@@
where $f_{\theta}(\mathbf{B}_{t})\in\mathbb{R}^{b\times|Y|}$ is the softmax output of the model. This criterion favors models whose predictions are simultaneously confident and diverse, reducing the risk of collapse when adapting to a novel domain (see alternative initialization in Appendix˜F).
@@ 5 Experiments
Baselines. We evaluate our method against several state‐of‐the‐art TTA baselines (for more details, see the Appendix˜G). For single-target TTA, we compare with TENT [39]. For continual TTA, we consider CoTTA [41] (using the affine-parameter variant, CoTTA*), RoTTA [44], ETA [26], and SAR [27]. To assess long-term stability, we include persistent TTA methods such as RDumb [30], PeTTA [12], EATA [26] (with Fisher-based regularization), and ROID [23] (reported as ROID*—a version omitting the augmentation consistency loss). For ViT-based models, we also compare with domain-disentanglement approaches like CoLA [3] and DPCore [46]. In segmentation experiments, we evaluate segmentation variants of TENT, CoTTA*, and BECoTTA [19]. For fair comparison, all methods update only the backbone’s affine parameters—except when we compare ReservoirTTA and BECoTTA (using LoRA) and DPcore (using visual prompts).
@@
Implementation Details. All methods are re-implemented in PyTorch [28] within a unified TTA repository [23] for fair comparison, using pre-trained source models from RobustBench [11]. See Appendix˜H for further implementation details.
@@ 5.1 Main Results
Classification. Table˜1 highlights the limitations of existing TTA methods when repeatedly adapted under recurring domain shifts and prolonged testing with three CNN-based backbones. (1) Single-target TTA: Methods like TENT initially adapt well but suffer from severe error accumulation over multiple cycles—for example, on CIFAR-10-C, error rises from 19.3% at visit 1 to 87.8% at visit 20, with similar trends on CIFAR-100-C and ImageNet-C. Plugging in ReservoirTTA significantly improves performance, though error accumulation still occurs, highlighting the need for both variance control modules from our theoretical analysis. (2) Continual TTA: Approaches like CoTTA and RoTTA mitigate short-term instability yet still accumulate errors with repeated adaptation. ReservoirTTA decouples domain-specific adaptation to prevent parameter drift—reducing, for instance, ETA’s CIFAR-10-C error from 30.9% to 16.4% at the 20th visit (Recurring CSC) and similarly improving ImageNet-C performance. (3) Persistent TTA: Although methods such as EATA and ROID are designed to prevent catastrophic forgetting, they exhibit limited re-adaptation (Figure˜1), with minimal improvements over repeated visits. ReservoirTTA overcomes this by enabling controlled adaptation without excessive re-learning—reducing EATA’s ImageNet-C error from 55.9% to 51.0% in Recurring CSC and ROID’s from 55.5% to 52.1%.
As shown in Table˜2, using a ViT-B-16 backbone on ImageNet-C further underscores ReservoirTTA’s advantages in recurrent continual TTA. Domain-disentangled methods like DPCore improve early adaptation but lack long-term stability. For instance, in Recurring CSC, CoLA reduces ETA’s 20th-visit error to 33.8%, whereas ReservoirTTA lowers it further to 31.9%. In CCC, CoLA stabilizes ETA at 40.2%, whereas ReservoirTTA achieves 38.5%. DPCore also struggles: in CCC, its error rises from 42.2% to 43.1%, whereas combining ReservoirTTA with the same prompt-tuning approach (VPT) [14]—used by DPCore for fair comparison—keeps the error at 39.2%. This gap shows that the domain identification mechanism in ReservoirTTA outperforms those in CoLA and DPCore by avoiding gradual degradation and accelerating performance improvement. As shown in Appendix˜D, our style features form higher-quality clusters than the ViT-based features used in CoLA and DPCore. Moreover, our model discovery mechanism is less sensitive to test batch order and estimates the number of domains more accurately (see Figures˜4 and I), resulting in fewer trainable parameters compared to DPCore and COLA. See Tables 1225 in Appendix˜K for full results across all 20 recurrences, CNN backbones (group norm), DomainNet-126, and PACS.
@@
Table 2: Average classification error (%) on ImageNet-C under recurring continual TTA (ViT-B/16). For CCC, we average over an adaptation window (e.g., steps 6701–40200). Means over 5 seeds. % Train Params = fraction of trainable parameters; Time = $\times$ vs. Source. Margins show negative/positive error changes vs. the base plug-in.
@@
Recurring visit Recurring visit Adaptation Step $\%$ Train Params
@@
[…]
@@
Figure 3: Sensitivity to reservoir size $K^{\text{max}}$ on CIFAR100-C (recurring CSC/CDC).